OpenStudy (anonymous):

The sum of two numbers is 70. The larger number is 4 more than twice the smaller number. What is the smaller number? NEED HELP PLEASE?

OpenStudy (sarahann):

The larger number is 48, the smaller number is 22. 48 is twice the amount of 22 plus 4. If the smaller number is "x" then the larger number is 2x+4. We know that x+2x+4 =70 which ius the same as saying 3x+4 = 70 or 3x = 70-4 = 66 so x, the smaller number is 66/3 =22

OpenStudy (anonymous):

thanks